OpenStudy (marmar10):

Help! Prove: 1-2sin^2 x/2sinxcosx = cot 2x

OpenStudy (marmar10):

\[\frac{ 1-2\sin ^{2} x}{ 2sinxcosx }= \cot 2x\]

OpenStudy (campbell_st):

well I think you'll find the denominator is sin(2x) = 2sin(x)cos(x) and the numerator \[\cos(2x) = \cos^2(x) - \sin^2(x)\] see if you can use a trig identity and substitute it to get the numerator... then it might become obvious.